𝑦(2π‘₯𝑦 + 1)𝑑π‘₯ βˆ’ π‘₯𝑑𝑦 = 0


1
Expert's answer
2021-08-10T15:15:36-0400

Write in the form of a first order Bernoulli Ordinary Differential Equation


"y'-\dfrac{1}{x}y=2y^2"

The general solution is obtained by substituting


"z=y^{1-n}=y^{1-2}=y^{-1}"

Differentiating, we find:


"z'=(y^{-1})'=-y^{-2}y'"

Solve


"y^{-2}y'-\dfrac{1}{x}y^{-1}=2"


"-z'-\dfrac{1}{x}z=2""z'+\dfrac{1}{x}z=-2"

"\mu(x)=e^{\int{1 \over x}dx}=e^{\ln x}=x"

We can make sure that the functionΒ "x" is the integrating factor


"xz'+z=-2x"

"(xz)'=-2x"

Integrate


"\int d(xz)=-\int2xdx"

"xz=-x^2+C"

"xy^{-1}=-x^2+C"

"y=\dfrac{x}{-x^2+C}"